OpenStudy (anonymous):

Identify the value of the x-intercept by estimating -b/m...y=2x-2 The y intercept is (0, -2)

OpenStudy (anonymous):

here m=2 and b=2 =-b/m=-(-2)/2=1

OpenStudy (anonymous):

got it